# Two Channel Multi impurity Kondo model: RKKY induced criticality

### v4

@@ Abstract
Introduction. The study of quantum impurity models has played a crucial role in the development of modern theories of strongly correlated electron systemsHewson (1993). Beginning with Anderson’s single impurity model and Wilson’s analysis of the renormalization group of the Kondo problem, these models have provided profound insights into the emergence of nontrivial many-body phenomena such as spin screening, quantum entanglement, and the breakdown of Fermi liquid theory. Theoretical advances in impurity models have not only led to fundamental concepts such as the Kondo effect, heavy-fermion behavior, and quantum criticality but have also been instrumental in approaching the many body problem on the lattice, as for example in dynamical mean-field theory (DMFT), which treats lattice problems by mapping them onto effective impurity models.
@@
Quantum impurity models are also important from the perspective of unconventional quantum critical points that are not well described by the Hertz-Millis-Moriya theory of quantum criticality. These zero temperature critical points are characterized by charge or spin susceptibilities that exhibit weak momentum dependence, while their frequency dependence follows a logarithmic form over a broad frequency range, and a specific heat behaving as $C(T)\sim T\ln T$ reminiscent of the two-channel Kondo impurity model. Salient examples include the marginal Fermi liquid theory describing optimally doped cuprates Varma et al. (1989) and the quantum critical point of heavy fermion systems for which the term local quantum criticality Si et al. (2001) is widely used. Interestingly, while the two-channel Kondo impurity model captures the logarithmic scaling of various observables over a broad range of energies, a more singular behavior emerges at much lower temperatures or frequencies in both heavy fermions Oeschler et al. (2008) and cuprates while still being conformally invariant as expected from a quantum impurity model Guo et al. (2024).
@@
The study of models with multiple impurities, brings new features arising from the RKKY interaction which have motivated experiments in mesoscopic systems as well as molecules absorbed on surfaces Craig et al. (2011); Bork et al. (2011); Spinelli et al. (2015). Keller et al. (2015). See ref Eickhoff and Anders (2021) for a recent context and a proper identification of the number of channels.
@@
The standard paradigm suggests that the competition between the RKKY interaction and the Kondo screening results in a quantum critical point when these two scales are comparable. This scenario has been explored by Jones and Varma Jones and Varma (1989), whose work was later generalized by Georges and Sengupta Georges and Sengupta (1995, 1994) to the two channel situation. The three impurity Kondo model, displays more a complex quantum critical point, and was studied in ref. Ingersent et al. (2005). In this letter we depart from this scenario and show that in the two channel Kondo model the RKKY interaction itself can generate new types of quantum criticality even when it is much smaller than the single impurity two channel Kondo scale. We solve exactly a model of the spin sector of the multichannel N impurity model, and show that it exhibits qualitatively new critical features when the number of impurities $N\geq 3$. We study a QIM (quantum impurity model), in the same universality class of the topological Kondo model studied in refs. Beri and Cooper (2012); Altland and Egger (2013); Altland et al. (2014a, b); Li et al. (2023). It exhibits a non trivial fixed point, arising directly from the presence of the RKKY interaction. The model exhibits two channel Kondo behavior at intermediate energies and crosses over to a more singular behavior at lower energies making it a possible candidate for the local quantum critical point governing the heavy fermions and the cuprates. This QIM has many other interesting potential applications and we illustrate this with some implications to the physics of magnetic impurities embedded in a Majorana metal and to heavy fermion materials, which host two channel Kondo lattice physics such as $UBe_{13}$ox (1994) and PrV2Al20 Sakai and Nakatsuji (2011).
@@
Construction of the effective theory. We start with the theory of a single spin coupled to two channels of conduction electrons in three dimensions via an isotropic spin interaction. It is well established that in this case a part of the bulk degrees of freedom decouple from the impurity such that the resulting effective Hamiltonian can be recast in terms of the three species of Majorana fermions. The Kondo temperature $T_{K}$ is exponential in $1/J$ and the effective Hamiltonian $E<<T_{K}$ is given by Coleman et al. (1995)
@@
At low energy, and to the leading order the low energy excitations are described by three non interacting Majorana fields $\chi^{a}$. $\xi$ is a drone variable, a local Majorana zero mode, which enables the representation of the impurity spin operator at energies $<<T_{K}$ as:
@@
However, when considering many impurities one should not forget that the fermions screening different impurities belong to the same D-dimensional bulk so that their correlation functions do not vanish. We will take this fact into account by integrating out part of the fermionic fields not located at the impurity sites. To this end we use Hamiltonian (1) and perform fusion of the four-fermion terms located at different sites:
@@
with $A_{2ch}$ is given by (3). The ultraviolet cut-off is $v/r$, where $r$ is an average distance between the impurities. The resulting model is a multichannel version of the model for the topological Kondo effect studied in Beri and Cooper (2012); Altland and Egger (2013); Altland et al. (2014a, b).
The multichannel version (9) was recently studied in Li et al. (2023), where it was established that it is critical. Therefore contrary to the conventional wisdom, the criticality is generated rather than destroyed by the RKKY interaction, at least provided the impurity concentration is sufficiently small so that the conditions (8) are fulfilled.
@@
a special case of the two-impurity models considered in Ingersent et al. (1992); Georges and Sengupta (1995) which included many couplings related to scattering between the impurities.
@@
The first term in the Hamiltonian (12) is irrelevant, but it prevents the model from being exactly solvable in the entire range of temperatures from $T_{K}$ to zero. The exact Bethe ansatz solution is possible for $|\bar{I}_{1}|=|\bar{I}_{2}|<\bar{I}_{3}$ only asymptotically at $T<<T_{K}$. The solution is similar to the one for spin $S$ XXZ Heisenberg Hamiltonian Kirillov and Reshetikhin (1987) and has been discussed in Kirillov and Reshetikhin (1985); Tsvelik (1995). We discus the results below, for more details on the the Bethe ansatz treatment see Sec.II of the Supplemental Material.
In contrast with the $N=2$ case the RKKY interaction is not marginal. It follows from (5) that the sign of the coupling is always positive and hence the interaction in (12) is always marginally relevant. Then its growth generates its own crossover scale $T_{RKKY}$. In the isotropic case we have $T_{RKKY}\sim T_{K}\exp[-1/2\rho(\epsilon_{F})\bar{I}]$. A trajectory goes from the single impurity two channel QCP SU2(2) at high energies to the multi-impurity 6-channel QCP at low energies since $O_{3}(3)\equiv SU_{6}(2)$, it is the 6-channel Kondo. The thermodynamics can be extracted from Tsvelik (1985) (see also Supplement). From the Bethe anzats solution the impurity contribution to the ground state entropy is given by $S(0)=\ln\sqrt{2+\sqrt{2}}.$ The leading irrelevant operator which determines the temperature dependence of the impurity contribution to the specific heat, is the first Kac-Moody descendant of the primary field in the adjoint representation $J_{-1}\Phi_{adj}$. Its conformal dimension at the critical point is $h=j(j+1)/(k+2)$ which for $j=1,k=6$ yields 1/4. The impurity free energy scales as $F_{imp}\sim T^{2h_{1}+1}$, so the specific heat is
@@
As we have stated above, model (9) is a multichannel version of the model describing the topological Kondo effect studied in Beri and Cooper (2012); Altland and Egger (2013); Altland et al. (2014a, b) and was recently studied in Li et al. (2023). $J_{ij}=\mbox{i}(\vec{\chi}_{i}\vec{\chi}_{j})$ are $O_{3}(N)$ Kac-Moody currents ($N$ is the number of impurities) and $s_{ij}=\mbox{i}\xi_{i}\xi_{j}$ are generators of the $o(N)$ algebra. Hence we have $n=3$ channel Kondo problem with orthogonal symmetry.
@@
The $O_{n}(N)$ sector absorbs the impurities and becomes quantum critical while the $O_{N}(n)$ sector of the bulk remains idle. The resulting critical theory is the $O_{n}(N)$ Wess-Zumino-Novikov-Witten (WZNW) model subject to a conformally invariant boundary condition. Its basic properties are described in Kimura (2021). The conformal dimensions of the primary fields of the $O_{n}(N)$ WZNW model are $h_{i}=\frac{C_{i}}{n+N-2},$ where $C_{i}$ are $O(N)$ group Casimirs in the corresponding representations.
@@
Magnetic impurities in a Majorana spin liquid with a Fermi surface The above results are directly applicable to a Majorana metal with magnetic impurities. Such metal is a spin liquid where elementary excitations are neutral Majorana fermions with a Fermi surface and Z2 gauge fluxes. This situation is realized for certain 3D latticees in an exactly soluble generalization of the famous Kitaev model of Z2 spin liquid Kitaev (2006) - the Yao-Lee (YL) model Hermanns and Trebst (2014); O’Brien et al. (2016); Coleman et al. (2022). We provide a more detailed description in the Supplementary Material. The 3D YL model has a phase transition below which the static Z2 gauge field can be ignored and the low energies excitations are three species of noninteracting Majorana fermions. The YL model does not have a charge sector. Therefore a coupling a cluster of impurity $S=1/2$ spins to the Yao Lee model with a Fermi surface provides a straightforward realization of the physics described in this paper.
@@
where $n_{i}$ is the impurity concentration and $\rho(\epsilon_{F})$ is the density of states at the Fermi surface. Recall that for the electrons it would be square root frequency dependence Affleck and Ludwig (1993). It enters in the expression for the thermal conductivity. Hence in the limit of small impurity concentration the YL model with spin impurities behaves as marginal FL with linear T dependence in the thermal resistivity. The low temperature behavior for the multi-impurity case is described above.
@@
with $\epsilon_{k}$ is the conduction electron dispersion $\sigma$ and $m$ spin and channel indices, running over two values, the index i runs over the sites of the lattice and $\mathbf{S}_{i}$ the spin at lattice site i. It was proposed as relevant to the heavy fermion materials such as UBe13, which exhibits a large resistivity down to its superconducting transition temperature $PrRh_{2}Zn_{20}$ which exhibits quadrupolar orderIshii et al. (2013). Not much is known about this model, even in limiting cases. Within single site DMFT which is exact in infinite dimensions, its normal phase realizes a non Fermi liquid incoherent metal and phases with broken symmetry such as channel orderingInui and Motome (2020) Hoshino et al. (2013). Slave particle studies reveal even more exotic properties Ge and Komijani (2022, 2024). In the paramagnetic phase, cluster and single site DMFT studies of this model do not exhibit Fermi liquid behavior.
Our fixed point is a candidate to describe this non Fermi liquid behavior. DMFT studies of this model in the paramagnetic phase. The approach maps the lattice problem into a cluster of N impurities spins embedded in a two channel bath obeying a DMFT self consistency condition, as summarized in the Supplementary Material. In cluster DMFT the impurity self energy, $\Sigma_{imp}$ is viewed as a functional of the hybridization function, and is obtained solving the impurity model. The correlation functions can be computed from the impurity model. Our discussion heavily relies on the assumption that the self energy of the type of impurity model does not diverge at low frequency. Then as shown in the Supplementary Material, the self consistent hybridization function is regular (i.e. it has a finite density of states) at low energies in a self consistent way. Having a regular hybridization function, the results of this paper apply and we expect a non Fermi liquid phase, starting with a regime of decoupled impurities at high temperatures, and flowing to the new fixed point identified in this paper throughout the paramagnetic phase.
@@ Conclusions.
Figure 2: The suggested picture of the renormalization group flow for the two channel Kondo model with multiple impurities. There are two critical points: the unstable Jones-Varma one and a new stable Wess-Zumino-Novikov-Witten critical point described in the paper.
@@
The scaling dimension of the spin operator at the low energy QCP is $1/(N+1)$ making its correlation function more singular as $N$ increases. At the same time the specific heat singularity and the impurity entropy per spin are reduced with the increase of $N$. This is consistent with the interpretation that the criticality is connected to freezing of the spin degrees of freedom due to the ordering or to the formation of some spin liquid state. In the context of cluster DMFT, the effective number of impurities is determined dynamically, with the effective range determined by the temperature starting from single impurity behavior at high temperaturesTanasković et al. (2011).
It would be interesting to relate the mechanism for criticality introduced in this paper, with the observations of criticality found in recent cluster DMFT studies of the Hubbard model at optimal doping Haule and Kotliar (2007); Sordi et al. (2011), Gleis et al. (2024) and to the experiments which exhibit two channel Kondo quantum criticality followed by a more singular rise of the response at low temperatures.
I Supplementary Material
@@ I.1 Magnetic impurities in a Majorana spin liquid with a Fermi surface
A Majorana metal is a spin liquid where elementary excitations are Majorana neutral fermions with a Fermi surface. This situation is realized in an exactly soluble spin model, which generalizes the famous Kitaev model of Z2 spin liquid Kitaev (2006) - the Yao-Lee (YL) model and has been investigated in several lattices Hermanns and Trebst (2014); O’Brien et al. (2016); Coleman et al. (2022).
@@
Since the commutation relations of the Majorana current operator coincide with the SU2(2) Kac-Moody current, the dynamics of the impurity spin is the same as in the 2-channel Kondo model (see Coleman et al. (1995)). The exact solution can be generalized on the case of anisotropic exchange interaction as, for example, in Tsvelik (1995). Alternatively, the solution or the U(1) symmetric case $J_{x}=J_{y}=J_{\perp}\neq J_{z}$ can be obtained by means of bosonization as in Georges and Sengupta (1994).
@@
The anisotropic Ising coupling between orbitals induces Majorana fractionalization Yao and Lee (2011) of spins and orbitals:
@@
Here, $\hat{u}_{ij}=ib^{\alpha_{ij}}_{i}b^{\alpha_{ij}}_{j}$ are the static $\mathbb{Z}_{2}$ gauge fields, (i.e. $[H_{YL},\hat{u}_{ij}]=0$).
@@
Figure 3: Majorana Fermi surface of the YL model on hyperoctagon lattice. Adopted from paper Coleman et al. (2022).
I.2 Bethe ansatz for $N=3$
I.3 T=0 results
The exact solution of model (12) of the main text is possible in the limit $T_{RKKY}/T_{K}\rightarrow 0$ when one can neglect the first term. In that limit model (12) with ¯$\textasciimacron\tilde{I}_{1}=\tilde{I}_{2}<\tilde{I}_{3}$ is equivalent to the XXZ 6-channel Kondo model. The Bethe ansatz equations for the XXZ 6-channel Kondo model are
@@ I.3 T=0 results
at $T=0$. The result will show that the power law dependence of ${\cal M}$ on ${\cal H}$ at ${\cal H}<T_{RKKY}$ is independent on the anisotropy.
@@
The integration limit $B$ is determined by the condition that the bulk magnetization is ${\cal M}_{bulk}={\cal H}/\Lambda$ or equivalently by the condition $\epsilon(B)=0$. Equations (26,27) are the Wiener-Hopf ones and can be solved analytically as, for example, in Tsvelick and Wiegmann (1985). The result for the magnetic moment is
@@
I.3.1 The thermodynamic equations
The TBA equations for the $k$-channel Kondo model with anisotropy and S=1/2 impurity are Tsvelik (1985):
@@ I.3.1 The thermodynamic equations
I.4 Multichannel Kondo lattice

### v5

@@ Abstract
Introduction. The study of quantum impurity models has played a crucial role in the development of modern theories of strongly correlated electron systems[1]. Beginning with Anderson’s single impurity model and Wilson’s analysis of the renormalization group of the Kondo problem, these models have provided profound insights into the emergence of nontrivial many-body phenomena such as spin screening, quantum entanglement, and the breakdown of Fermi liquid theory. Theoretical advances in impurity models have not only led to fundamental concepts such as the Kondo effect, heavy-fermion behavior, and quantum criticality but have also been instrumental in approaching the many body problem on the lattice, as for example in dynamical mean-field theory (DMFT), which treats lattice problems by mapping them onto effective impurity models.
@@
Quantum impurity models are also important from the perspective of unconventional quantum critical points that are not well described by the Hertz-Millis-Moriya theory of quantum criticality. These zero temperature critical points are characterized by charge or spin susceptibilities that exhibit weak momentum dependence, while their frequency dependence follows a logarithmic form over a broad frequency range, and a specific heat behaving as $C(T)\sim T\ln T$ reminiscent of the two-channel Kondo impurity model. Salient examples include the marginal Fermi liquid theory describing optimally doped cuprates [2] and the quantum critical point of heavy fermion systems for which the term local quantum criticality [3] is widely used. Interestingly, while the two-channel Kondo impurity model captures the logarithmic scaling of various observables over a broad range of energies, a more singular behavior emerges at much lower temperatures or frequencies in both heavy fermions [4] and cuprates while still being conformally invariant as expected from a quantum impurity model [5].
@@
The study of models with multiple impurities, brings new features arising from the RKKY interaction which have motivated experiments in mesoscopic systems as well as molecules absorbed on surfaces [6, 7, 8]. [9]. See Ref. [10] for a recent context and a proper identification of the number of channels.
@@
The standard paradigm suggests that the competition between the RKKY interaction and the Kondo screening results in a quantum critical point when these two scales are comparable. This scenario has been explored by Jones and Varma [11], whose work was later generalized by Georges and Sengupta [12, 13] to the two channel situation. The three impurity Kondo model, displays more a complex quantum critical point, and was studied in Ref. [14]. In this letter we depart from this scenario and show that in the two channel Kondo model the RKKY interaction itself can generate new types of quantum criticality even when it is much smaller than the single impurity two channel Kondo scale. We solve exactly a model of the spin sector of the multichannel N impurity model, and show that it exhibits qualitatively new critical features when the number of impurities $N\geq 3$. We study a QIM (quantum impurity model), in the same universality class of the topological Kondo model studied in refs. [15, 16, 17, 18, 19]. It exhibits a non trivial fixed point, arising directly from the presence of the RKKY interaction. The model exhibits two channel Kondo behavior at intermediate energies and crosses over to a more singular behavior at lower energies making it a possible candidate for the local quantum critical point governing the heavy fermions and the cuprates. This QIM has many other interesting potential applications and we illustrate this with some implications to the physics of magnetic impurities embedded in a Majorana metal and to heavy fermion materials, which host two channel Kondo lattice physics such as UBe13[20] and PrV2Al20 [21],[22].
@@
Construction of the effective theory. We start with the theory of a single spin coupled to two channels of conduction electrons in three dimensions via an isotropic spin interaction. It is well established that when the channels is coupled symmetrically a part of the bulk degrees of freedom decouple from the impurity such that the resulting effective Hamiltonian can be recast in terms of the three species of Majorana fermions. The two channel Kondo effect becomes quantum critical. The stability of the critical point may be achieved for magnetic ions with non-Kramers ground state such as Pr ones [21]. The Kondo temperature $T_{K}$ is exponential in $1/J$ and the effective Hamiltonian $E<<T_{K}$ is given by [23]
@@
At low energy, and to the leading order the low energy excitations are described by three non interacting Majorana fields $\chi^{a}$. $\xi$ is a drone variable, a local Majorana zero mode, which enables the representation of the impurity spin operator at energies $T_{K}$ as:
@@
However, when considering many impurities one should not forget that the fermions screening different impurities belong to the same D dimensional bulk so that their correlation functions do not vanish. We will take this fact into account by integrating out part of the fermionic fields not located at the impurity sites. To this end we use Hamiltonian (1) and perform fusion of the four-fermion terms located at different sites:
@@
with $A_{2ch}$ is given by (3). The ultraviolet cut-off is $v/r$, where $r$ is an average distance between the impurities. The resulting model is a multichannel version of the model for the topological Kondo effect studied in [15, 16, 17, 18].
The multichannel version (9) was recently studied in [19], where it was established that it is critical. Therefore contrary to the conventional wisdom, the criticality is generated rather than destroyed by the RKKY interaction, at least provided the impurity concentration is sufficiently small so that the conditions (8) are fulfilled.
@@
a special case of the two-impurity models considered in [24, 12] which included many couplings related to scattering between the impurities.
@@
The first term in the Hamiltonian (12) is irrelevant, but it prevents the model from being exactly solvable in the entire range of temperatures from $T_{K}$ to zero. The exact Bethe ansatz solution is possible for $|\bar{I}_{1}|=|\bar{I}_{2}|<\bar{I}_{3}$ only asymptotically at $T<<T_{K}$. The solution is similar to the one for spin $S$ XXZ Heisenberg Hamiltonian [25] and has been discussed in [26, 27]. We discus the results below, for more details on the the Bethe ansatz treatment see Sec.II of the Supplemental Material.
In contrast with the $N=2$ case the RKKY interaction is not marginal. It follows from (5) that the sign of the coupling is always positive and hence the interaction in (12) is always marginally relevant. The coupling constants flow under RG as is it customary for non-Abelian theories. Its growth generates its own crossover scale $T_{RKKY}$. In the isotropic case we have $T_{RKKY}\sim T_{K}\exp[-1/2\rho(\epsilon_{F})\bar{I}]$. A trajectory goes from the single impurity two channel QCP SU2(2) at high energies to the multi-impurity 6-channel QCP at low energies since $O_{3}(3)\equiv SU_{6}(2)$, it is the 6-channel Kondo. The thermodynamics can be extracted from [28] (see also Supplement). From the Bethe anzats solution the impurity contribution to the ground state entropy is given by $S(0)=\ln\sqrt{2+\sqrt{2}}.$ The leading irrelevant operator which determines the temperature dependence of the impurity contribution to the specific heat, is the first Kac-Moody descendant of the primary field in the adjoint representation $J_{-1}\Phi_{adj}$. Its conformal dimension at the critical point is $h=j(j+1)/(k+2)$ which for $j=1,k=6$ yields 1/4. The impurity free energy scales as $F_{imp}\sim T^{2h_{1}+1}$, so the specific heat is
@@
As we have stated above, model (9) is a multichannel version of the model describing the topological Kondo effect studied in [15, 16, 17, 18] and was recently studied in [19]. $J_{ij}=\mbox{i}(\vec{\chi}_{i}\vec{\chi}_{j})$ are $O_{3}(N)$ Kac-Moody currents ($N$ is the number of impurities) and $s_{ij}=\mbox{i}\xi_{i}\xi_{j}$ are generators of the $o(N)$ algebra. Hence we have $n=3$ channel Kondo problem with orthogonal symmetry.
@@
The $O_{n}(N)$ sector absorbs the impurities and becomes quantum critical while the $O_{N}(n)$ sector of the bulk remains idle. The resulting critical theory is the $O_{n}(N)$ Wess-Zumino-Novikov-Witten (WZNW) model subject to a conformally invariant boundary condition. Its basic properties are described in [29]. The conformal dimensions of the primary fields of the $O_{n}(N)$ WZNW model are $h_{i}=\frac{C_{i}}{n+N-2},$ where $C_{i}$ are $O(N)$ group Casimirs in the corresponding representations.
@@
Magnetic impurities in a Majorana spin liquid with a Fermi surface The above results are directly applicable to a Majorana metal with magnetic impurities. Such metal is a spin liquid where elementary excitations are neutral Majorana fermions with a Fermi surface and Z2 gauge fluxes. This situation is realized for certain 3D latticees in an exactly soluble generalization of the famous Kitaev model of Z2 spin liquid [30] - the Yao-Lee (YL) model [31, 32, 33]. We provide a more detailed description in the Supplementary Material. The 3D YL model has a phase transition below which the static Z2 gauge field can be ignored and the low energies excitations are three species of noninteracting Majorana fermions. The YL model does not have a charge sector. Therefore a coupling a cluster of impurity $S=1/2$ spins to the Yao Lee model with a Fermi surface provides a straightforward realization of the physics described in this paper.
@@
where $n_{i}$ is the impurity concentration and $\rho(\epsilon_{F})$ is the density of states at the Fermi surface. Recall that for the electrons it would be square root frequency dependence [34]. It enters in the expression for the thermal conductivity. Hence in the limit of small impurity concentration the YL model with spin impurities behaves as marginal FL with linear T dependence in the thermal resistivity. The low temperature behavior for the multi-impurity case is described above.
@@
with $\epsilon_{k}$ is the conduction electron dispersion $\sigma$ and $m$ spin and channel indices, running over two values, the index i runs over the sites of the lattice and $\mathbf{S}_{i}$ the spin at lattice site i. It was proposed as relevant to the heavy fermion materials such as UBe13, which exhibits a large resistivity down to its superconducting transition temperature, PrRh2Zn20 which exhibits quadrupolar order[35] and PrV2Al20 [21]. Not much is known about this model, even in limiting cases. Within single site DMFT which is exact in infinite dimensions, its normal phase realizes a non Fermi liquid incoherent metal and phases with broken symmetry such as channel ordering[36, 37, 21]. Slave particle studies reveal even more exotic properties [38, 21]. In the paramagnetic phase, cluster and single site DMFT studies of this model do not exhibit Fermi liquid behavior.
Our fixed point is a candidate to describe this non Fermi liquid behavior. DMFT studies this model in the paramagnetic phase. The approach maps the lattice problem into a cluster of N impurities spins embedded in a two channel bath obeying a DMFT self consistency condition, as summarized in the Appendix. In cluster DMFT the impurity self energy, $\Sigma_{imp}$ is viewed as a functional of the hybridization function, and is obtained solving the impurity model. The correlation functions can be computed from the impurity model. Our discussion heavily relies on the assumption that the self energy of the type of impurity model does not diverge at low frequency. Then as shown in the Appendix, the self consistent hybridization function is regular (i.e. it has a finite density of states) at low energies in a self consistent way. Having a regular hybridization function, the results of this paper apply and we expect a non Fermi liquid phase, starting with a regime of decoupled impurities at high temperatures, and flowing to the new fixed point identified in this paper throughout the paramagnetic phase.
@@ Conclusions.
Figure 2: The suggested picture of the renormalization group flow for the two channel Kondo model with multiple impurities. There are two critical points: the unstable Jones-Varma one and a new stable Wess-Zumino-Novikov-Witten critical point described in the paper. The phase on the right from the JV-point is magnetic.
@@
The scaling dimension of the spin operator at the low energy QCP is $1/(N+1)$ making its correlation function more singular as $N$ increases. At the same time the specific heat singularity and the impurity entropy per spin are reduced with the increase of $N$. This is consistent with the interpretation that the criticality is connected to freezing of the spin degrees of freedom due to the ordering or to the formation of some spin liquid state. In the context of cluster DMFT, the effective number of impurities is determined dynamically, with the effective range determined by the temperature starting from single impurity behavior at high temperatures[39].
It would be interesting to relate the mechanism for criticality introduced in this paper, with the observations of criticality found in recent cluster DMFT studies of the Hubbard model at optimal doping [40, 41], [42] and to the experiments which exhibit two channel Kondo quantum criticality followed by a more singular rise of the response at low temperatures.
I Appendix
@@ I.1 Magnetic impurities in a Majorana spin liquid with a Fermi surface
A Majorana metal is a spin liquid where elementary excitations are Majorana neutral fermions with a Fermi surface. This situation is realized in an exactly soluble spin model, which generalizes the famous Kitaev model of Z2 spin liquid [30] - the Yao-Lee (YL) model and has been investigated in several lattices [31, 32, 33].
@@
Since the commutation relations of the Majorana current operator coincide with the SU2(2) Kac-Moody current, the dynamics of the impurity spin is the same as in the 2-channel Kondo model (see [23]). The exact solution can be generalized on the case of anisotropic exchange interaction as, for example, in [27]. Alternatively, the solution or the U(1) symmetric case $J_{x}=J_{y}=J_{\perp}\neq J_{z}$ can be obtained by means of bosonization as in [13].
@@
The anisotropic Ising coupling between orbitals induces Majorana fractionalization [43] of spins and orbitals:
@@
Here, $\hat{u}_{ij}=ib^{\alpha_{ij}}_{i}b^{\alpha_{ij}}_{j}$ are the static $\mathbb{Z}_{2}$ gauge fields, (i.e. $H_{YL},\hat{u}_{ij}]=0$).
@@
Figure 3: Majorana Fermi surface of the YL model on hyperoctagon lattice. Adopted from paper [33].
I.2 Bethe ansatz for $N=3$. T=0 results
The exact solution of model (12) of the main text is possible in the limit $T_{RKKY}/T_{K}\rightarrow 0$ when one can neglect the first term. In that limit model (12) with $\textasciimacron\tilde{I}_{1}=\tilde{I}_{2}<\tilde{I}_{3}$ is equivalent to the XXZ 6-channel Kondo model. The Bethe ansatz equations for the XXZ 6-channel Kondo model are
@@
[…]
@@
The integration limit $B$ is determined by the condition that the bulk magnetization is ${\cal M}_{bulk}={\cal H}/\Lambda$ or equivalently by the condition $\epsilon(B)=0$. Equations (26,27) are the Wiener-Hopf ones and can be solved analytically as, for example, in [44]. The result for the magnetic moment is
@@
I.2.1 Bethe ansatz for $N=3$. The thermodynamic equations
The TBA equations for the $k$-channel Kondo model with anisotropy and S=1/2 impurity are [28]:
@@
I.3 Multichannel Kondo lattice